OpenStudy (anonymous):

find the implicit derivative of -2x^3+(x^2)y-3y^3=-1

OpenStudy (amistre64):

its the same rules as for explicit

OpenStudy (amistre64):

"but not when it has a product of y" is a misstatement. the product rule still applies to terms that are a product of y

OpenStudy (anonymous):

ok sure, thanks amistre64, I know how to solve it but may have difficulty explaining it, so I will leave it to you

OpenStudy (amistre64):

\[D[-2x^3+x^2y-3y^3=-1]\] \[D[-2x^3]+D[x^2y]-D[3y^3]=D[-1]\] exponent rule, product rule, exponent rule(and chain rule for the y), constant rule

OpenStudy (anonymous):

would the final answer be \[6x ^{2}/(2x-9y ^{2})\]

OpenStudy (amistre64):

hmm, quite possible. i like the 6x^2 aprt; and the division looks good to

OpenStudy (amistre64):

but let me check \[D[-2x^3]+D[x^2y]-D[3y^3]=D[-1]\] \[-6x^2+D[x^2]y]+x^2D[y]-9y^2y'=0\] \[-6x^2+2xy+x^2y'-9y^2y'=0\] a little off

OpenStudy (amistre64):

that chain rule seems to have been the culprit in yours \[-6x^2+2xy+x^2y'-9y^2y'=0\] \[x^2y'-9y^2y'={6x^2-2xy}\] \[y'(x^2-9y^2)={6x^2-2xy}\] \[y'=\frac{6x^2-2xy}{x^2-9y^2}\]